# Accounting for Focus Ambiguity in Visual Questions

### v1

@@ Abstract
No existing work on visual question answering explicitly accounts for ambiguity regarding where the content described in the question is located in the image. To fill this gap, we introduce VQ-FocusAmbiguity, the first VQA dataset that visually grounds each region described in the question that is necessary to arrive at the answer. We then provide analysis showing how our dataset for visually grounding ‘questions’ is distinct from visually grounding ‘answers’, and characterize the properties of the questions and segmentations provided in our dataset. Finally, we benchmark modern models for two novel tasks: recognizing whether a visual question has focus ambiguity and localizing all plausible focus regions within the image. Results show that the dataset is challenging for modern models. To facilitate future progress on these tasks, we publicly-share the dataset with an evaluation server at https://focusambiguity.github.io/.
@@ 1 Introduction
Ambiguous language is a common part of communication. It entails using vague words or phrases that can be interpreted in multiple plausible ways, ideally alongside context clarifying the intended meaning. For example, when a three-year-old asks “What is this?”, we can understand the meaning if the child simultaneously points to an item (e.g., a red pomegranate). Despite the widespread use of ambiguous language in visual questions, to our knowledge, no prior AI work accounts for under-specified language in questions grounded to multiple regions in the images. This concept is called focus ambiguity [9], and additional examples are shown in Figure 1.
Figure 1: Visual questions with and without focus ambiguity (aka. question ambiguity). Focus Ambiguity means the content described in the questions (underlined words) refers to multiple plausible regions in the images. Focus groundings (solid lines with regions filled) can be different from answer groundings [4] (dotted lines). For each example, the same color is used for each segmentation overlaid on the image and its associated text.
To fill this gap, we introduce the first dataset challenge centered on accounting for focus ambiguity in visual questions. For improved readability, throughout this paper, we refer to “focus ambiguity in visual questions” as “ambiguous questions”. Our dataset, called VQ-FocusAmbiguity, consists of 4,357 examples where we segment (aka, ground) all plausible image regions to which the content described in each question could refer. We focus only on the content described in each question that is essential to reason to the answer. Our dataset provides a nearly even distribution between examples containing versus lacking question ambiguity. We next characterize the dataset. Crucially, as exemplified in Figure 1-(3)(5) and Figure 6-(b), we found that when question ambiguity happens, the segmentations for the questions are usually different from segmentations showing where answers are grounded in the images. Finally, we benchmark modern models and found they struggle when repurposed for our two novel tasks of (1) recognizing whether a visual question has focus ambiguity and (2) locating all plausible image regions to which the content described in the question could refer. We publicly-share the dataset and an evaluation server to facilitate future progress on developing AI models that address these challenging tasks: https://focusambiguity.github.io/.
@@
Success in this work can immediately benefit today’s users of visual question answering (VQA) services, spanning blind and sighted individuals, who already regularly ask visual questions using mobile phone apps (e.g., Be My AI, Microsoft’s Seeing AI), smart glasses (e.g., Meta’s Ray Bans, Envision AI), and the web (e.g., Q&A platforms such as Stack Exchange). It would enable AI agents to alert users when there is question ambiguity as well as interactively guide users towards disambiguating their intent by having them specify which from all plausible image regions the question is referring to. This work can also support enhanced reasoning abilities of AI agents, by encouraging an intermediate step of determining a question’s focus in an image towards deciding what answer to predict. Finally, we expect this work will serve as a pioneering example for addressing focus ambiguity for other vision-language tasks (e.g., image captioning and visual storytelling), in languages beyond English (e.g., by translating our dataset), and across more modalities (e.g., focus ambiguity for questions asking about visual, audible, tactile, or olfactory content).
@@ 2 Related Work
Ambiguity in Question Answering.
Automated systems account for several types of ambiguity in question answering. Most focus on a purely language setting [36, 27, 13, 22], such as when a word has multiple plausible meanings (e.g., “revolting” can mean rebelling or disgusting). A few explore the multimodal VQA setting [37, 29, 35, 2, 40, 15], by accounting for when and why different answers are observed from different people (e.g., due to subjectivity and differing levels of granularity). Our work complements prior work by being the first to disentangle the questions themselves as a source of ambiguity, thereby helping clarify how different answers can arise by showing the reasoning process from a question that leads to the answers.
@@
Many datasets exist for VQA [34, 18, 26, 5, 1, 11] and locating the visual evidence showing where answers to visual questions reside [10, 23, 20, 38, 3, 4]. We instead introduce the first dataset that locates where the content described in the question could be referring to within an image.
@@
More generally, this work will contribute to literature on locating linguistic expressions in an image, whether with coarse detections (i.e., rectangles) or fine-grained segmentations (i.e., detailed outlines). Examples include object detection [42], instance segmentation [14, 19], referring expression comprehension [30, 6], and described object detection [33, 39]. Our work instead centers on localizing a visual question’s content, which yields a diversity of linguistic expressions including vague terms such as “this”, “it”, and “her” (e.g., third person pronouns, singular demonstrative pronouns).
3 Dataset
We now describe how we created the VQ-FocusAmbiguity dataset and then we characterize its properties.
@@ 3.1 Dataset Creation
Each example in our dataset has three parts: an image, a question, and segmentations for all regions that could be the focus of the question. We created the dataset by extending four diverse sources, that are described in Table 1. The visual data spans content that (1) shows a single object and complex scenes; (2) comes from sighted and visually impaired photographers, and (3) has objects at various locations, of many sizes. Questions ask about many subjects—including about objects and their parts—as well as about their relationships and actions, using vague terms (e.g., “this”), specific categories (e.g., “bus”), and detailed descriptions (e.g., “person in blue next to the car”).
Dataset Image Source Question Source Ambiguity Labels Segmentations % Unambiguous (#)
PACO-LVIS [31] COCO (2017) [14, 24] Synthesized + Workers Workers Workers* 50% (1,700)
@@
Workers
Visually Impaired People
@@
Table 1: Description of the four data sources used in VQ-FocusAmbiguity. Entries in bold represent new annotations created by our team. (* denotes when annotators selected between candidate segmentations rather than creating them from scratch)
@@ 3.1.1 Extensions of Segmentation Datasets
Most examples are derived from two-entity segmentation datasets, which already provide images with segmentations.
@@
We extend PACO-LVIS [31] and MSRA-B [25]. PACO-LVIS is built on the COCO images [24], which were collected from online image-sharing platforms (e.g., Flickr) and contain at least one of 80 object categories that would be easily recognizable by a four-year-old within complex scenes showing two or more objects. PACO provides exhaustive segmentations for all instances belonging to 75 object categories and 200 part categories. MSRA-B, in contrast, is a salient object segmentation dataset designed to contain a single foreground object per image, without consideration of the object category [21]. While images from MSRA-B are similarly scraped from online image-sharing platforms, they only show a single salient object. Segmentations of all the salient objects are provided.
@@
To mitigate the likelihood that AI models would be trained on data included in our dataset, we leveraged only the test sets of both sourcese randomly sampled 2,501,700mples from PACO-LVIS. For MSRA-B images, we focused on the subset of images that prior work indicated show “a single, noncontroversial foreground object of interest” [17] and then removed all examples containing human faces and adult content (determined using the GPT-4o model [28]). This left us with 627 images.
@@
For the segmented images from the data sources, we next established questions paired with relevant segmentations that could be the focus of the content described in the questions.
For PACO-LVIS, we achieved this using a home-grown annotation interface that showed the image with all available segmentations overlaid on the image. The interface first prompted the annotator to generate a question by presenting a list of AI-suggested questions [^1] Due to space constraints, we described how the questions were generated by AI in the Supplementary Materials and letting the annotator choose between either (1) authoring a question from scratch, (2) selecting a suggested candidate question as is, or (3) selecting a suggested candidate question after refining it. [^2] When generating ambiguous questions, annotators chose options 1, 2, and 3 for 55%, 31%, and 14% respectively. For unambiguous questions, annotators chose options 1, 2, and 3 for 43%, 43%, and 14% respectively. Next, the user was prompted to select all segmentations to which the question could be grounded.
For MSRA-B, we only generated unambiguous questions and used the single segmentation associated with each image as the question’s focus. We automatically generated a question for each image, using variants of the most common question from authentic visual question asked by people with vision impairments [18], ”What is this?”. We also introduced further diversity with questions like ”What is prominent?” and ”What is in the foreground?”.
@@ 3.1.2 Extensions of Visual Question Answering Datasets
The remaining examples are built on two VQA datasets.
@@
We extend VizWiz-VQA [18] and VQAv2 [12]. VizWiz-VQA represents an authentic use case where people with visual impairments asked questions about images they took. VQAv2 is the most popular VQA dataset for model benchmarking and was created by asking people to make up questions about images that would “stump a robot”. We focused on the 4,440 examples from these sources contained in the publicly-available splits of the VQA-AnswerTherapy dataset [4]; i.e., its training and validation sets, to enable comparison of question groundings and answer groundings. In particular, VQA-AnswerTherapy provides segmentations showing where each of the valid answer resides.
@@
In-house annotators labeled every example from the data source by (1) locating the phrase in the question that needed to be grounded to the image in order to answer the question and (2) indicating whether there was ambiguity regarding where the phrase is referring to in the image. This culminated in 165 ambiguous questions and 4,275 unambiguous questions, with 85% (i.e., 3792/4440) belonging to visual questions with a single answer grounding and 15% (i.e., 648/4440) belonging to visual questions with multiple answer groundings. The annotators then generated segmentations for all 165 ambiguous questions and 165 randomly sampled unambiguous questions, locating all regions the relevant phrase could focus on in the images.
@@ 3.1.3 Dataset Splits
This culminated in 4,357 examples with 4,357 classification labels and 9,685 instance segmentation. [^3] To facilitate future research, we also publicly-share the metadata about the additional 4,110 examples from the VQA datasets that we flagged as unambiguous but did not segment. Examples are nearly evenly distributed between containing and lacking focus ambiguity, with 1,865 and 2,492 respectively. Given the relatively small size of the dataset, we design it to support a few-shot learning setting. For both the training and validation sets, we randomly sample 10 unambiguous and 10 ambiguous examples from each data source – for unambiguous questions, we have four sources (PACO-LVIS, MSRA-B, AnswerTherapy-VQAv2, and AnswerTherapy-VQA), and for ambiguous questions, we have three sources (PACO-LVIS, AnswerTherapy-VQAv2, and AnswerTherapy-VQA). This resulted in training, validation, and test splits with 70 (40 unambiguous and 30 ambiguous), 70 (40 unambiguous and 30 ambiguous), and 4,217 (1,805 unambiguous and 2,412 ambiguous) examples respectively.
@@ 3.2 Dataset Analysis
erize how questions compare for those with versus without focus ambiguity. We provide analysis with respect to the entire VQ-FocusAmbiguity dataset as well as with respect to each data source.
Figure 2: Summative statistics characterizing the number of words in questions for ambiguous versus unambiguous questions overall as well as with respect to each data source. For each box, the line in the middle denotes the median length, the edges denote the 25th and 75th percentile lengths, and the whiskers extend to the most extreme lengths. The hollow circles are outliers.
Statistics regarding how many words are in the questions are shown in Figure 2. Overall, we observe a tendency for unambiguous questions to contain more words, particularly for examples from PACO-LVIS. We hypothesize this correlation stems from more the additional words providing extra context for disambiguating the intended questions.
@@
Figure 3: Icicle chart showing the first three words for all questions (a) with focus ambiguity and (b) without focus ambiguity. Each rectangle size is proportional to the number of questions with that word/phrase, with the leftmost column showing the first word and each subsequent column showing a subsequent word.
We next characterize common linguistic patterns in questions by showing how the distribution of the first three words compares for questions with versus without focus ambiguity. We demonstrate this with the icicle chart shown in Figure 3. There are similarities across both types of questions, including that most commonly begin with “What is”. A notable distinction though is that questions with focus ambiguity more often begin with “Is the”, whereas questions without focus ambiguity more often start with “How many”. These highlight a contrast where ambiguous questions focus on a single entity while unambiguous questions permit focusing on multiple entities, since “Is” grammatically should be followed by a singular word/phrase while “How many” permits acknowledging more than one entity. Intuitively, it makes sense that ambiguity is more likely to arise when a question is framed in a singular form rather than a plural form, since the former does not permit acknowledging multiple image regions (e.g., “What color is the kite?” versus “What color are the kites?”) We hypothesize a valuable predictive cue for AI models deciding whether there is focus ambiguity is to consider whether a question asks about a single versus multiple entities.
@@
To further investigate our hypothesis, we utilized NLTK to flag whether any word in each question contains a plural noun (i.e., a plural common noun or plural proper noun). Supporting our hypothesis, we found that questions lacking focus ambiguity are more than three times as likely to contain a plural noun than questions with focus ambiguity, with 21.1% versus 4.52% respectively. This trend was more pronounced for PACO and for VQAv2, where questions were typed, than for VizWiz-VQA where the questions were initially spoken and subsequently transcribed. [^4] Specifically, 4.2% of ambiguous questions contained plural nouns while 29% of unambiguous questions for PACO-LVIS, versus 4.2% and 17.6% respectively for VQA-Therapy (7.7% and 26.8% for VQA-v2; 0% and 8% for VizWiz-VQA). Despite this slight difference in trends, the presence of plural nouns is not sufficient alone to determine whether there is focus ambiguity.
@@
We report the position of instance segmentations by computing the center of mass of each instance segmentation with respect to the entire image. Each coordinate can range from 0 to 1, and we normalize all images to ensure (x,y) values are comparable to each other. Results are shown in Figure 4. There is a greater bias for instance segmentations lacking focus ambiguity to be located in the center of an image, due to the salient object data source MSRA-B. However, we observe that questions lacking focus ambiguity can also manifest the diverse locations typical for questions containing focus ambiguity, particularly when examining the unambiguous examples created for the more complex images coming from PACO-LVIS.
@@
Summative statistics regarding how many instance segmentations are associated with each question are shown in Figure 5a. We exclude unambiguous examples since, by definition, they only contain one segmentation. We observe similar statistics across all data sources, with an overall median of 3 and mean of 3.97 segmentations per question.
We next measure the fraction of image pixels occupied by each instance segmentation. Results are shown in Figure 5b. Examples lacking ambiguity tend to have segmentations occupying a larger portion of the image, likely because the tendency in such cases is for images to feature a single, dominant salient object. However, we do observe unambiguous examples with very small coverage like that observed for ambiguous questions, particularly for the PACO-LVIS data source. Consequently, properties of an instance’s image coverage alone is insufficient for predicting whether there is question ambiguity.
Additionally, we analyze the prevalence of objects versus parts for serving as the instance segmentations for the PACO-LVIS data source. Within PACO, 75.3% of instance segmentations are of only objects, 22.4% are of only parts, and 2.3% feature a mix of objects and parts.
@@
Question Groundings Versus Answer Groundings.
For the 330 examples sourced from the VQA-AnswerTherapy dataset, we flagged all examples whether the instance segmentations in our focus ambiguity dataset match answer groundings. [^5] This labeling process is described in the Supplementary Materials. We observed different trends for the different types of questions. For the ambiguous questions, 79% (131 out of 165) had groundings that are different for the question and answers. In contrast, 64% (106 out of 165) of questions lacking focus ambiguity had groundings that are matching for the questions and answers. Examples of both scenarios are shown in Figure 6. These findings help motivate the importance of locating a question’s focus as an important, independent stepping stone towards providing end users of VQA services all valid answers.
@@
For 265 examples with focus ambiguity, we manually coded the reasons for focus ambiguity. We used all 91 examples from VQAv2, 74 from VizWiz, and a random sample of 100 from PACO. We identified two primary reasons:
@@
  • Multiple instances of the same singular category account for 61.5% overall, with 84.6% (i.e., 77) in VQAv2, 4% (i.e., 3) in VizWiz-VQA, and 83% (i.e., 83) in PACO. An example is “What is next to the mirror?” when multiple mirrors are present.
Other rare reasons include (a) perspective ambiguity (e.g., “Who is pulling the other side?”), (b) subjectivity (e.g., “What is the most distinctive feature on the building?”), (c) language ambiguity (e.g., “What is in the picture?” while it can either refer to the entire image or a painting in the image), (d) multiple plausible parts for the same entity (e.g., “What is the part of the lamp that is fully visible?”).

### v2

# Acknowledging Focus Ambiguity in Visual Questions

@@ Abstract
No published work on visual question answering (VQA) accounts for ambiguity regarding where the content described in the question is located in the image. To fill this gap, we introduce VQ-FocusAmbiguity, the first VQA dataset that visually grounds each plausible image region a question could refer to when arriving at valid answers. We next analyze and compare our dataset to existing datasets to reveal its unique properties. Finally, we benchmark modern models for two novel tasks related to acknowledging focus ambiguity: recognizing whether a visual question has focus ambiguity and locating all plausible focus regions within the image. Results show that the dataset is challenging for modern models. To facilitate future progress on these tasks, we publicly share the dataset with an evaluation server at https://vizwiz.org/tasks-and-datasets/focus-ambiguity-in-visual-questions/.
@@ 1 Introduction
Ambiguous language is a common part of communication. It entails using vague words or phrases that can be interpreted in multiple plausible ways, ideally alongside context clarifying the intended meaning. For example, when a three-year-old asks “What is this?”, we can understand the meaning if the child simultaneously points to an item (e.g., a red pomegranate). However, context is not always provided to clarify the intended meaning of a question, as exemplified in Figure 1. In such cases, the language in the question could refer to multiple plausible visual regions. We call this focus ambiguity in visual questions, or more concisely “ambiguous questions” and “question ambiguity”.
Our paper is motivated by the belief that a VQA system should notify users when there is question ambiguity and then facilitate them to arrive at the desired interpretation. The possible repercussions from VQA services not providing such information can be grave, potentially inflicting adverse social, professional, legal, financial, and personal consequences. For instance, imagine if the answer to the question in Figure 1 about “What is the cleaning product?” leads a blind person to use window cleaner to wash their dishes they will use for eating instead of the dish soap. Alternatively, imagine replacing the question in Figure 1 with “What is the medicine?” when instead three pill bottles are visible. The current obstacle for developing ambiguity-aware VQA solutions is that no benchmark dataset exists for establishing to what extent VQA models are aware of question ambiguity and so where further improvement is needed.
Figure 1: Visual questions with focus ambiguity, meaning language in the questions (underlined words/phrases) refer to multiple plausible image regions (aka, segmentations). ‘Q’ denotes the question, ‘F’ denotes the focus specified in the question, and ‘A’ denotes the answer. We created a new dataset that originates from four sources to represent ambiguous and unambiguous questions, with two of the sources originating from existing VQA datasets [15, 22] and exemplified in this figure. As shown, the location of the question grounding and answer grounding can match (in which case, matching colors are used for all segmentations and their associated text) and can differ (bottom right example, where an additional answer segmentation is shown with a dashed line).
To fill this gap, we introduce the first dataset goal-oriented towards focus ambiguity in visual questions. Called VQ-FocusAmbiguity, it consists of 5,500 examples and segments (aka, grounds) all plausible image regions to which the language in each question could refer. The dataset has a nearly even distribution between ambiguous and unambiguous examples. We also characterize this dataset and how it relates to a related answer grounding dataset. Crucially, as exemplified in Figure 1, our analysis underscores the importance of disentangling questions as a source of ambiguity, since the focus of language in questions can differ from visual evidence showing answers. Additional scenarios where the question and answer groundings can differ include for questions about relations (e.g., “What is above the mirror?”, with the question grounding of the mirror, and answer grounding of objects above the mirror) and activities (e.g., “What is the person doing?”, with the question grounding of a person and the answer grounding of a frisbee). Finally, we show that modern models perform poorly for the tasks of (1) recognizing whether a visual question has focus ambiguity and (2) locating all plausible image regions to which the content described in the question could refer. Our fine-grained analysis reveals where models struggle, and so provides valuable insights for future efforts in model development.
@@
Success in developing ambiguity-aware solutions can immediately benefit today’s users of VQA services, spanning blind and sighted individuals, who already regularly ask visual questions using mobile phone apps (e.g., Be My AI, Microsoft’s Seeing AI), smart glasses (e.g., Meta’s Ray Bans, Envision AI), and the web (e.g., Q&A platforms such as Stack Exchange). It would enable AI agents to alert users when there is question ambiguity as well as interactively guide users towards disambiguating their intent by having them specify which from all plausible image regions the question is referring to. This work can also support enhanced reasoning abilities of AI agents, by encouraging an intermediate step of determining a question’s focus in an image towards deciding what answer to predict. Finally, we expect this work will serve as a pioneering example for addressing focus ambiguity for other vision-language tasks (e.g., image captioning [11], visual storytelling [34], language-guided image editing [37]), and across more modalities (e.g., focus ambiguity for questions asking about visual, audible, tactile, or olfactory content).
@@ 2 Related Work
Automatically Acknowledging Ambiguity in Question Answering.
Automated systems already account for several types of ambiguity in question answering. Most focus on a purely language setting [49, 38, 17, 27], such as when a word has multiple plausible meanings (e.g., “revolting” can mean rebelling or disgusting). A few focus on multimodal VQA [50, 40, 46, 4, 53, 19], by accounting for when and why different answers are observed from different people (e.g., due to subjectivity and differing levels of granularity). Our work complements prior work by being the first to disentangle the questions themselves as a source of ambiguity, thereby helping clarify how different answers can arise by showing the reasoning process from a question that leads to the answers.
@@
Many datasets exist for VQA [45, 22, 36, 7, 2, 15] and locating the visual evidence showing where answers to visual questions reside [55, 14, 29, 24, 51, 5, 6]. One dataset, Visual7W [55], even locates where the language in questions refers to in images. Extending prior work, we introduce the first dataset goal-oriented to focus ambiguity, with examples showing when language in questions refers to multiple plausible image regions.
@@
More generally, this work contributes to existing literature on locating linguistic expressions in an image, which is already explored for tasks like object detection [56], instance segmentation [18, 23], referring expression comprehension [41, 8], and described object detection [44, 52]. Complementing these tasks, our work locates where language in a question refers to in an image, which yields diverse linguistic expressions, including vague terms such as “this”, “it”, and “her” (e.g., third person pronouns, singular demonstrative pronouns).
3 VQ-FocusAmbiguity Dataset
@@ 3.1 Dataset Creation
Each example in our dataset has three parts: an image, a question, and segmentations for all regions that could be the focus of the question. We created the dataset by extending four diverse sources, that are described in Table 1. The visual data spans content that (1) shows a single object and complex scenes; (2) comes from sighted and visually impaired photographers, and (3) has objects at various locations and of many sizes. Questions ask about many subjects—including about objects and their parts—as well as about their relationships and actions, using vague terms (e.g., “this”), specific categories (e.g., “bus”), and detailed descriptions (e.g., “person in blue next to the car”).
Dataset Image Source Question Source Ambiguity Labels Segmentations % Unambig (#)
PACO [42] COCO (2017) [18, 31] Synthesized + Workers Workers Workers* 50% (2,272)
@@
Workers [31]
Visually Impaired People [22]
@@
Table 1: Description of the four data sources used in VQ-FocusAmbiguity. Entries in bold represent new annotations created by our team. (* denotes when annotators selected between candidate segmentations rather than creating them from scratch; Unambig = Unambiguous).
@@ 3.1.1 Extensions of Segmentation Datasets
Most examples are derived from two entity segmentation datasets, which already provide images with segmentations.
@@
We leverage test sets of PACO [42] and MSRA-B [33], which both contain images scraped from online image-sharing platforms. PACO uses the COCO images [31], which each show a complex scene with two or more objects. PACO provides exhaustive segmentations for all instances belonging to 75 object and 200 part categories. MSRA-B, in contrast, is a salient object segmentation dataset designed to contain a single foreground object per image, agnostic to the object category [26].
@@
For PACO, we randomly sampled 2,272 examples. For MSRA-B, we focused on the 626 images that both lacked human faces and adult content (we determined this using the GPT-4o model [39]) and were flagged as containing “a single, noncontroversial foreground object of interest” [21].
@@
We next established questions with language referring to the segmentations.
For PACO, we achieved this using a home-grown annotation interface that showed the image with all available segmentations overlaid on the image. The interface first prompted the annotator to generate a question by presenting a list of AI-suggested questions and letting the annotator choose between either (1) authoring a question from scratch, (2) selecting a suggested candidate question as is, or (3) selecting a suggested candidate question after refining it. [^1] When generating ambiguous questions, annotators chose options 1, 2, and 3 for 55%, 31%, and 14% respectively. For unambiguous questions, annotators chose options 1, 2, and 3 for 43%, 43%, and 14% respectively. Next, the user was prompted to select all segmentations to which the question could be grounded.
For MSRA-B, we only generated unambiguous questions and used the single segmentation per image as the question’s focus. We automatically generated the questions, using variants of the most common question asked by people with vision impairments [22]: “What is this?”
@@ 3.1.2 Extensions of Visual Question Answering Datasets
The remaining examples extend two VQA datasets.
@@
We extend VizWiz-VQA [22] and VQAv2 [16]. VizWiz-VQA represents an authentic use case where people with visual impairments asked questions about images they took. VQAv2 is the most popular VQA dataset for model benchmarking and was created by asking people to make up questions about images that would “stump a robot”. We focused on the 4,440 examples from these sources contained in the publicly-available splits of the AnswerTherapy dataset [6] (i.e., its training and validation sets) to enable comparison between its answer groundings and our question groundings.
@@
We established a trusted annotation protocol through three iterative refinement steps, discussed in the supplementary materials. Following this protocol, in-house annotators labeled every example from the data source by (1) locating the phrase in the question that needed to be grounded to the image in order to answer the question and (2) indicating whether there was ambiguity where the phrase is referring to in the image. This culminated in 165 ambiguous questions and 4,275 unambiguous questions, with 85% (i.e., 3792/4440) belonging to visual questions with a single answer grounding and 15% (i.e., 648/4440) belonging to visual questions with multiple answer groundings. The annotators then segmented all 165 ambiguous questions and 165 randomly sampled unambiguous questions, locating all regions the relevant phrase could focus on in the images.
Data annotation culminated in 5,500 visual questions with 12,880 instance segmentations and 5,500 classification labels (the binary flag for ambiguity is inferred from the number of segmentations). [^2] To facilitate future research, we also publicly-share the metadata about the additional 4,110 examples from the VQA datasets that we flagged as unambiguous but did not segment. Examples are nearly evenly distributed between containing and lacking question ambiguity, with 2,437 and 3,063 examples respectively.
@@ 3.1.3 Dataset Splits
Consistent with recent published VQA datasets [54, 7, 30, 35, 25, 47], we split this dataset to support zero/few-shot learning settings. This recent trend reflects that state-of-the-art performance regularly arises from foundation models in such settings; e.g., Frozen in 2021 [48], Flamingo in 2022 [1], ViTiS in 2023 [13], and LLaVA-v1.6 in 2024 [32]. For both the training and validation sets, we randomly sampled 10 unambiguous and 10 ambiguous examples from each data source to increase domain diversity. With four sources for unambiguous questions (PACO, MSRA-B, AnswerTherapy-VQAv2, and AnswerTherapy-VizWiz) and three sources for ambiguous questions (PACO, AnswerTherapy-VQAv2, and AnswerTherapy-VizWiz), we end up with 70, 70, and 5,360 examples in our training, validation, and test splits respectively.
@@ 3.2 Dataset Analysis
We first characterize how questions compare for examples with versus without focus ambiguity. We provide analysis with respect to the entire VQ-FocusAmbiguity dataset as well as with respect to each data source.
@@
Figure 2: Analysis of questions in the dataset. (a) Box plot showing the number of words in questions that are unambiguous (UA) versus ambiguous (A), overall and for each data source (outliers are omitted for improved readability). (b) Icicle chart showing the first three words for all questions with and without question ambiguity. Each rectangle size is proportional to the number of questions with that word/phrase, with the left column showing the first word and each subsequent column showing a subsequent word.
Statistics regarding how many words are in the questions are shown in Figure 2(a). Overall, we observe a tendency for unambiguous questions to contain more words, particularly for examples from PACO. We hypothesize this correlation stems from extra words providing additional context that disambiguates the intended questions. Unambiguous questions also exhibit greater variation in length, as reflected by a higher standard deviation of 3.13 for unambiguous questions versus 1.97 for ambiguous questions.
@@
We next characterize common linguistic patterns in questions by visualizing the distribution of their first three words in Figure 2(b). The key distinction between both question types is that questions with focus ambiguity more often begin with “Is the” while questions without focus ambiguity more often start with “How many”. Intuitively, it makes sense that ambiguity is more likely to arise when a question is framed in a singular form (“Is a”) rather than a plural form (“How many”), since the former does not permit acknowledging multiple image regions (e.g., “What color is the kite?” versus “What color are the kites?”). To further investigate this intuition, we utilized NLTK to flag whether any word in each question contains a plural noun (i.e., a plural common noun or plural proper noun). Supporting our hypothesis, we found that unambiguous questions are more than three times as likely to contain a plural noun than ambiguous questions, with 23.8% versus 4.7% respectively. This trend was more pronounced for PACO and for VQAv2, where questions were typed, than for VizWiz-VQA where the questions were initially spoken and subsequently transcribed. [^3] For PACO, plural nouns were found for 4.7% of ambiguous questions and 30.9% for unambiguous questions versus 4.2% and 17.6% for VQA-Therapy (7.7% and 26.8% for VQA-v2; 0% and 8% for VizWiz-VQA). Despite this slight difference in trends, the presence of plural nouns is not sufficient alone to determine whether there is question ambiguity.
@@
We report the position of instance segmentations by computing the center of mass of each instance segmentation with respect to the entire image. Each coordinate can range from 0 to 1, and we normalize all images to ensure (x,y) values are comparable to each other. Results are shown in Figure 3. There is a greater bias for instance segmentations lacking focus ambiguity to be located in the center of an image, likely due to the salient object data source MSRA-B. However, we observe that unambiguous questions can also manifest the diverse locations typical for ambiguous questions, particularly for the unambiguous examples coming from the more complex images in PACO.
@@
Summative statistics regarding how many instance segmentations are associated with each question are shown in Figure 4(a). We exclude unambiguous examples since, by definition, they contain one segmentation. We observe similar statistics across all data sources, with an overall median of 3 and mean of 4 segmentations per ambiguous question.
We next measure the fraction of image pixels occupied by each instance segmentation. Results are shown in Figure 4(b). Unambiguous questions tend to have segmentations occupying a larger portion of the image, likely because the tendency in such cases is for images to feature a single, dominant salient object. However, we do observe unambiguous examples with very small coverage like that observed for ambiguous questions, particularly for the PACO data source. Consequently, properties of an instance’s image coverage alone is insufficient for predicting whether there is question ambiguity.
Additionally, we analyze the prevalence of objects versus parts for serving as the instance segmentations for the PACO data source. Within PACO, 81.4% of instance segmentations are of only objects, 15.8% are of only parts, and 2.8% feature a mix of objects and parts.
Question Groundings Versus Answer Groundings.
We flagged for all examples from the AnswerTherapy dataset (i.e., 330) whether the instance segmentations in our question groundings match the answer groundings. We observed different trends for the different types of questions. For the ambiguous questions, 79% (131 out of 165) had groundings that are different for the question and answers. In contrast, 64% (106 out of 165) of unambiguous questions had groundings that are matching for the questions and answers. Examples of both scenarios are shown in Figure 5. These findings reinforce the importance of locating a question’s focus as an important, independent stepping stone towards providing users of VQA services all valid answers.
@@
For 265 examples with question ambiguity, we manually coded the reasons for question ambiguity. We used all 91 examples from VQAv2, 74 from VizWiz, and a random sample of 100 from PACO. We identified two primary reasons:
  • Multiple instances of the same category account for 61.5% overall, with 84.6% (i.e., 77) in VQAv2, 4% (i.e., 3) in VizWiz-VQA, and 83% (i.e., 83) in PACO. An example is “What is next to the mirror?” when multiple mirrors are present.
@@
Other rare reasons include (a) perspective ambiguity (e.g., “Who is pulling the other side?”), (b) subjectivity (e.g., “What is the most distinctive feature on the building?”), (c) language ambiguity (e.g., “What is in the picture?” while it can either refer to the entire image or a painting in the image), and (d) multiple plausible parts for the same entity (e.g., “What is the part of the lamp that is fully visible?”).